# Multiple Streams of Relation Extraction: Enriching and Recalling in Transformers

### v1

@@ Abstract
When an LLM learns a relation during finetuning (e.g., new movie releases, corporate mergers, etc.), where does this information go? Is it extracted when the model processes an entity, recalled just-in-time before a prediction, or are there multiple separate heuristics? Existing localization approaches (e.g. activation patching) are ill-suited for this analysis because they tend to replace parts of the residual stream, potentially deleting information. To fill this gap, we propose dynamic weight-grafting between fine-tuned and pre-trained language models to show that fine-tuned language models both (1) extract relation information learned during finetuning while processing entities and (2) “recall" this information in later layers while generating predictions. In some cases, models need both of these pathways to correctly generate finetuned information while, in other cases, a single “enrichment" or “recall" pathway alone is sufficient. We examine the necessity and sufficiency of these information pathways, examining what layers they occur at, how much redundancy they exhibit, and which model components are involved—finding that the “recall" pathway occurs via both task-specific attention mechanisms and a relation extraction step in the output of the attention and the feedforward networks at the final layers before next token prediction.
@@ 1 Introduction
Large Language Models (LLMs) are capable of storing and recalling a large number of relationships and associations (Petroni et al., 2019), but what happens when we finetune pretrained LLMs to learn new relationships? How does a model encode this information in its parameters and what mechanisms extract this new information during text generation?
A line of interpretability work has focused on understanding how Transformer-based language models extract relation information by examining information flow through networks (Geva et al., 2023; Ferrando and Voita, 2024), directly editing model parameters (Meng et al., 2022), or by searching for interpretable “relation directions" in a Transformer’s residual stream (Hernandez et al., 2023). However, a central question remains: when we add new relationship information to an LLM, is that information added just to the entity (e.g., just in the embeddings) or is it localized more in response to the entity in higher layers closer to next token prediction?
Language models are updated with new factual information via finetuning all the time—presidents are elected, Popes are selected, and streaming services keep commissioning new movies—can we isolate the mechanism by which these new relationships are added to a Transformer LLM?
Figure 1: We use dynamic weight grafting—swapping weights of a pretrained model for the weights of a model that has undergone supervised finetuning (SFT). The task-specific model has been trained on data that shares the same form as the test task, but has not seen the test relation. The relation-specific model has been trained on the test relations. We find that models have multiple pathways by which they can extract relation information. One of these pathways uses both task-specific attention mechanisms on the first entity and the final token as well as relation extraction mechanisms in feed-forward networks at the final layers before next token prediction.
@@
Previous approaches to localizing relation extraction have either used variants of activation patching (Meng et al., 2022) or ablations (Geva et al., 2023) to see which components contribute to the ultimate next token prediction. However, activation patching and ablations have a key limitation: they operate by modifying or replacing activations inside the model, which unintentionally deletes the computations that came before. For example, when an activation at a specific layer and token position is patched or ablated, we also overwrite all the upstream information that was flowing into that activation. This makes it difficult to tell whether a component of the model is actively extracting new information, or simply passing along information that was computed earlier. As a result, it’s hard to isolate which mechanisms are truly responsible for incorporating finetuned relation knowledge.
@@
To address this, we propose dynamic weight grafting  [^1] Code available at: https://github.com/toddnief/knowledge-grafting, a method for studying how finetuned knowledge is used in language models by swapping in weights from a finetuned model during generation. These swaps can be done selectively—at specific layers, components, and token positions (see Figure 3(b))—allowing us to test which parts of the model are sufficient to reproduce the effects of finetuning, without disrupting the rest of the computation. This combines the advantages of model grafting, which leaves previous computations intact (Panigrahi et al., 2023; Ilharco et al., 2022) with the ability to apply causal mediation analysis to specific mechanisms in the model, similar to activation patching (Heimersheim and Nanda, 2024; Goldowsky-Dill et al., 2023; Meng et al., 2022).
Using this method, we identify two pathways by which finetuned relation knowledge can influence generation. In some cases, we confirm that models enrich entity representations early in the sequence and carry that information forward (Geva et al., 2023, 2022). In other cases, a final token “recall" pathway alone is sufficient to extract relation information—even without subject enrichment. This indicates that the later layers of a finetuned model contain a mechanism to recall information from finetuning, even in response to a representation that does not contain information from the finetuning set.
2 Background
Relation Extraction
@@ 2 Background
We focus on the relation extraction task, which has been studied extensively in the NLP literature (Zhao et al., 2024) and discussed in classic representation learning literature (Hinton et al., 1986). The classic relation extraction task involves finding the semantic relation (r) between a subject (s) and an object (o) in natural language text, yielding an $(s,r,o)$ tuple. In our setting, however, we focus on generative models, seeking to understand the mechanisms by which models correctly generate an object from an $(s,r,o)$ tuple given the subject and relation in a natural language prompt.
@@
Given some pre-trained and finetuned model parameters $\theta^{\text{pre}}$ and $\theta^{\text{ft}}$, how might one localize the mechanisms responsible for the change in model behavior in the finetuned model? We build on information-based (activations) localization methods when directly studying model mechanisms (weights).
Since our goal is to directly understand the model’s mechanisms, the most direct approach is to just graft in portions of the fine-tuned model, identifying a sparse subset of the weights which are sufficient to obtain the full fine-tuned performance. For instance, Panigrahi et al. (2023) define a grafted model $\boldsymbol{\tilde{\theta}}$ using a mask $\gamma$:
@@
where $i$ refers to the $i^{\text{th}}$ parameter of each model, “pre” refers to the pretrained model, and “ft” refers to the finetuned model. Equivalently, Ilharco et al. (2022) express it as
$\tilde{\boldsymbol{\theta}}=\boldsymbol{\theta}^{\text{pre}}+\boldsymbol{\gamma}\circ\left(\boldsymbol{\theta}^{\text{ft}}-\boldsymbol{\theta}^{\text{pre}}\right)$
Note however that this does not account for the possibility of mechanisms which activate only on certain tokens; furthermore, it only provides a notion of sufficiency, when we would also like to assess the necessity of these edits.
Causal Mediation Analysis via Activation Patching
To address these limitations, there has been a host of work which instead attempts localization based on the information flow through the residual stream vectors $\lambda(t,l)$ at token $t$ and layer $l$. The most popular of these approaches is broadly termed activation patching, which we use to refer to any method which replaces some vectors $\lambda(t,l)$ with new vectors $\tilde{\lambda}(t,l)$ (Heimersheim and Nanda, 2024; Goldowsky-Dill et al., 2023; Meng et al., 2022)
The primary advantage of activation patching is that it easily fits within a causal mediation analysis (Mueller et al., 2024; Vig et al., 2020). Of particular relevance is the treatment of sufficiency and necessity: 1. Sufficiency is achieved via a high natural indirect effect, e.g. “Does replacing the $\lambda^{A}(t,l)$ with $\lambda^{B}(t,l)$ cause the patched model to behave like model $B$?” 2. Necessity is achieved via a low natural direct effect, e.g. “Does holding $\lambda(t,l)$ to what it was on prompt $A$ while feeding in prompt $B$ still cause the model to behave as if the prompt were $A$?”
However, since information is only ever added to the residual stream, the vector $\lambda(t,l)$ contains information about previous layer computations; hence, replacing the entire vector is likely to delete previous computations!
Partial Information Replacement via Directional Patching
What if, rather than replacing the entire residual stream vector, we instead replaced only a portion of it? This is the idea behind directional patching (Wu et al., 2023; Tigges et al., 2023; Geiger et al., 2024)–namely to replace only the portion corresponding to some subspace $v$:
$\tilde{\lambda}(t,l)=\lambda^{A}-\mathrm{proj}_{v}\lambda^{A}+\mathrm{proj}_{v}\lambda^{B}$ (1)
However, typical methods for finding $v$ (e.g. mean difference between activations on different prompts, gradient descent, etc.) do not restrict to the current layer’s computations. This is fine if the goal is to trace information flow through the residual stream, but it is not restrictive enough to pinpoint which operations are responsible for the change in model behavior. If we wish to isolate which mechanisms are responsible for the change in model behavior, we need to restrict directional patching to only affect the current layer’s computations.
3 Dynamic Weight Grafting
We take the position that, at least for the relation extraction task, the most natural vector to patch to the residual stream at a given component is the one which model $B$ (e.g. a fine-tuned model) would have computed if given model $A$’s input (the base model). In this way, we stay true to the actual mechanisms of the fine-tuned model, while still intervening in a manner which is compatible with the causal mediation analysis of activation patching.
@@ 3 Dynamic Weight Grafting
That is, given two models $\boldsymbol{\theta}^{A}$ and $\boldsymbol{\theta}^{B}$, consider the ordered sequence of their weight matrices $\boldsymbol{\theta}^{A}=[\boldsymbol{\theta_{1}}^{A}\ldots\boldsymbol{\theta_{M}}^{A}]$ and $\boldsymbol{\theta}^{B}=[\boldsymbol{\theta_{1}}^{A}\ldots\boldsymbol{\theta_{M}}^{B}]$. Let $\boldsymbol{\gamma}$ be a $1\times M$ mask over these components. That is, each $\boldsymbol{\theta_{c}}^{A}$ corresponds to a specfic model component (e.g. the $W^{Q}$ matrix at the 12th layer, the up-projection matrix for the feedforward network at the 5th layer)  [^2] We could also consider each layer as a component, which we do in some of the experiments.. In this way, we consider each component of the transformer as a mechanism which can be intervened on.
We define Dynamic Weight Grafting as token-wise, component-wise weight grafting:
$\boldsymbol{\tilde{\theta}}_{m}(t)=\begin{cases}\boldsymbol{\theta}_{c}^{\text{A}}&\text{if }\gamma_{c}(t)=0\\ \boldsymbol{\theta}_{c}^{\text{B}}&\text{if }\gamma_{c}(t)=1\end{cases}$
@@
where $c$ refers to the $c^{\text{th}}$ component of each model and $t$ refes to the token position. This is illustrated in Figure 3(b). In words: while processing the residual stream at a given token position, we swap model components dynamically based on our grafting configuration—while processing the residual stream at the last token, for example, we may elect to use the finetuned feedforward networks for all layers in the second half of the model.
Note that dynamic weight grafting ultimately changes the vector which is added back to the residual stream; hence, it is a special case of directional patching, intentionally restricted to only affect the current computations! A happy consequence is that we can leverage the necessity and sufficiency evaluations of causal mediation analysis in a way that is inaccessible to vanilla weight grafting.
In most of our experiments, we consider $\theta_{i}^{\text{A}}$ to be the pretrained model $\theta_{i}^{\text{PRE}}$ and we consider $\theta_{i}^{\text{B}}$ to refer to the finetuned model $\theta_{i}^{\text{SFT}}$. Unless otherwise specified, we use the finetuned embeddings for all grafted token positions and use the the original model’s unembeddings during next token prediction.
@@ 4 Experiments & Results
We use four pretrained Transformer-based decoder-only language models in our experiments: Llama3 Grattafiori et al. (2024), Pythia 2.8b Biderman et al. (2023), GPT2-XL Radford et al. (2019), Gemma Team et al. (2024). Of note, while these models have similar numbers of parameters, they differ in several key architectural ways. See Table 2 in Section A.3 for a comparison of models.
@@
We follow Allen-Zhu and Li (2023) and use templated supervised finetuning data to control what relationship information models are exposed to during finetuning. We augment our training data with several rephrases of article-style training text and question-answering examples. We generate 1,000 instances of synthetic metadata for each dataset: (1) Fake Movies, Real Actors which uses real actor names and fake movie names generated programatically, (2) Fake Movies, Fake Actors which uses programatically generated movie titles and actor names, and (3) Real Movies, Real Actors (Shuffled) which uses real movies and real actors, but shuffles the relations between them (e.g. “Keanu Reeves starts in The Departed alongside Meryl Streep”). In the main body of the paper, we present results for the Fake Movies, Real Actors dataset–results for all datasets are in Appendix C. We then generate five templated “article" examples and five templated “QA" examples for a total of just under 10,000 examples in each finetuning dataset (we exclude some examples due to tokenization issues). See Section A.1.2 and Section A.1.3 for examples of templates and training examples. All models are trained using next token prediction. See Section A.3.3 for training details.
@@
Table 1: Relation prompt templates used to test model relation extraction capabilities
@@
Headline {first_actor} {relation} {relation_preposition} in a movie {preposition}
@@
QA Q: Who {relation} {relation_preposition} a movie {preposition} {first_actor}? A: An actor named
4.1 Which positions are sufficient for relation extraction?
Where is relation information activated? Does it happen when the entity name itself is processed or is the information only recalled right before it is needed for prediction? A priori, it’s not clear if model features build on each other in a way that each step of the extraction pipeline is necessary to extract a relation or if a handful of features each independently make a certain prediction more likely.

### v2

# Multiple Streams of Knowledge Retrieval: Enriching and Recalling in Transformers

@@ Abstract
When an LLM learns a new fact during finetuning (e.g., new movie releases, newly elected pope, etc.), where does this information go? Are entities enriched with relation information, or do models recall information just-in-time before a prediction? Or, are “all of the above” true with LLMs implementing multiple redundant heuristics? Existing localization approaches (e.g., activation patching) are ill-suited for this analysis because they usually replace parts of the residual stream, thus overriding previous information. To fill this gap, we propose dynamic weight grafting, a technique that selectively grafts weights from a finetuned model onto a pretrained model. Using this technique, we show two separate pathways for retrieving finetuned relation information: 1) “enriching" the residual stream with relation information while processing the tokens that correspond to an entity (e.g., “Zendaya” in “Zendaya co-starred with John David Washington”) and 2) “recalling" this information at the final token position before generating a target fact. In some cases, models need information from both of these pathways to correctly generate finetuned facts while, in other cases, either the “enrichment" or “recall" pathway alone is sufficient. We localize the “recall” pathway to model components—finding that “recall" occurs via both task-specific attention mechanisms and an entity-specific extraction step in the feedforward networks of the final layers before the target prediction. By targeting model components and parameters, as opposed to just activations, we are able to understand the mechanisms by which finetuned knowledge is retrieved during generation.
@@ 1 Introduction
When a new pope is elected and and we want an LLM to answer “Who’s the pope?” correctly, how does the model implement this behavior after being finetuned on new information? Large Language Models (LLMs) are capable of storing and retrieving a large number of relationships and associations from pretraining (Petroni et al., 2019; Roberts et al., 2020), but how does a model encode this information in its parameters? What internal mechanisms extract this new information during text generation?
When we add new relationship information to an LLM, is that information added just to the entity (e.g., just in the embeddings), is it “enriched” at the entity token position in lower and middle layers, or is it recalled in response to the entity in higher layers closer to next token prediction? And, on a more granular level, which model components contribute to remembering newly learned relation information?
(a) Dynamic weight grafting vs. Activation patching
(b) Position grafting vs. Component grafting
Figure 1: We introduce dynamic weight grafting—swapping the weights of a pretrained model for the weights of a model that has undergone supervised finetuning (SFT). (a) We compare dynamic weight grafting to activation patching. In activation patching, we replace model activations at a specific point with activations from another run. In dynamic weight grafting, we replace mechanism by swapping in specific parameter matrices of a finetuned model into a pretrained model. (b) A schematic showing the different dynamic weight grafting schemes used in our experiments. In position grafting, we use either the entire pretrained or finetuned model at a given token position. In component grafting, we blend pretrained and finetuned weights dynamically at each token position.
@@
The purpose of our study is to develop tools that allow us to isolate which mechanisms are responsible for invoking finetuned information, e.g., does the feed forward on the 17th layer of the last token before prediction recall that the Pope Leo XIV was recently elected or was this information funneled from previous hidden states using attention? Previous approaches to localizing relation extraction have either used variants of activation patching (Meng et al., 2022a) or ablations (Geva et al., 2023) to see which components contribute to next token prediction. Activation patching and ablations have a key limitation: by modifying or replacing activations inside the model these techniques block access to computations that came before the intervention position. For example, when an activation at a specific layer and token position is patched or ablated, we also overwrite all the upstream information that was flowing into that activation. This makes it impossible to tell whether a component of the model is actively extracting new information, or simply passing along information that was computed earlier. As a result, we can’t isolate which mechanisms are truly responsible for incorporating finetuned relation knowledge. Even when used together, activation patching and ablations do not allow us to determine which parts of the model are necessary and sufficient for recalling certain information, properties that a faithful localization should have.
@@
To address this, we propose dynamic weight grafting,  [^1] Code is available at https://github.com/toddnief/knowledge-grafting a method for studying how finetuned knowledge is used in language models by swapping in weights from a finetuned model during generation. These swaps can be done selectively—at specific layers, components, and token positions (see Figure˜1(b))—allowing us to test which parts of the model are necessary and sufficient to reproduce the effects of finetuning without disrupting the rest of the computation. This combines the advantages of model grafting, which leaves previous computations intact (Panigrahi et al., 2023; Ilharco et al., 2022) with the ability to apply causal mediation analysis to specific mechanisms in the model, similar to activation patching (Heimersheim and Nanda, 2024; Goldowsky-Dill et al., 2023; Meng et al., 2022a).  [^2] We also note that this method can be applied to any setting where we have a finetuned model and a pretrained model, including most post-training procedures.
Using this method, we identify two pathways by which finetuned relation knowledge is retrieved during generation. We see that, in some cases, grafting only at the first entity or only at the last token is sufficient to recover good relation completion performance. This implies two things: (1) If an entity is “enriched" with relation information, generic mechanisms can extract the correct entity to complete the relation tuple (Meng et al., 2022a; Geva et al., 2023; 2022) and (2) “recall" mechanisms can extract relation completions from entities that were not enriched with finetuned relation information. Both pathways together nearly recover full finetuning performance, and grafting everything except these pathways results in near-zero relation completion accuracy. In short, dynamic weight grafting allows us to show that the enrichment and recall pathways are necessary and sufficient for relation completion, localizing finetuned information with granularity that was not previously possible.
2 Methodology
Given pretrained and finetuned model parameters $\theta^{\text{pre}}$ and $\theta^{\text{ft}}$, how might one localize the mechanisms responsible for model behavior changes in the finetuned model? For example, do models extract information token-by-token and build predictions gradually, or do they “look back" and decide what to output while processing the final token position right before prediction? Historically, researchers have used activation patching to understand model behavior by intervening on the information flowing through the model. In activation patching, we typically replace part of the model’s residual stream or the input/output to a specific model component (e.g. attention block on layer 15). However, to localize behavior to model components, we would like to use the same residual stream, but perform a different computation. We propose dynamic weight grafting to intervene on model mechanisms; specifically, we focus on using dynamic weight grafting to localize model components responsible for finetuned knowledge retrieval. See Figure˜1 for a comparison between activation patching and weight grafting. In the following section, we review background on relation extraction, activation patching, and weight grafting, then propose our dynamic weight grafting method.
Relation Completion
@@ 2 Methodology
The classic relation extraction task involves finding the semantic relation (r) between a subject (s) and an object (o) in natural language text, yielding an $(s,r,o)$ tuple (Zhao et al., 2024; Hinton et al., 1986). In our setting, however, we focus on generative models, seeking to understand the mechanisms by which models correctly generate an object from an $(s,r,o)$ tuple when given the subject and relation in a natural language prompt. See Figure˜1(b) for an example.
Activation Patching
There has been a host of work which attempts localization based on the information flow through the residual stream vectors $\lambda(t,l)$ at token $t$ and layer $l$. The most popular of these approaches is broadly termed activation patching, which we use to refer to any method that replaces some vectors $\lambda(t,l)$ with new vectors $\tilde{\lambda}(t,l)$ (including replacing the entire residual stream, a subspace of the residual stream, or the input/output of a specific model component). (Heimersheim and Nanda, 2024; Goldowsky-Dill et al., 2023; Meng et al., 2022a).
Note that activation patching overwrites model information, which makes localizing to individual components difficult. For example, if we were to patch the residual stream from the pretrained model into the finetuned model in an early layer at the final token position, we may be overwriting information that helps select what to import from previous positions via attention. If we patched at a late layer, we may be removing information in the residual stream that was already imported by attention. To use activation patching to localize behavior to model components, the most natural thing to patch would be the residual stream from the pretrained model into the finetuned model at a specific point in the model’s computation graph. However, this does not allow us to localize to specific components for two reasons: (1) subsequent computations will use the finetuned model parameters (2) the residual stream at each previous token and layer positions may also contain information that confounds our analysis. To handle this, we instead focus on grafting in finetuned model parameters so that we can isolate model mechanism.
@@
In previous work, Panigrahi et al. (2023) define a grafted model $\bm{\tilde{\theta}}$ using a mask $\gamma$; they learn a sparse mask over model parameters to recover finetuned performance:
@@
where $i$ refers to the $i^{\text{th}}$ parameter of each model, “pre” refers to the pretrained model, and “ft” refers to the finetuned model. Equivalently, Ilharco et al. (2022) express this idea as: $\tilde{\bm{\theta}}=\bm{\theta}^{\text{pre}}+\bm{\gamma}\circ\left(\bm{\theta}^{\text{ft}}-\bm{\theta}^{\text{pre}}\right)$
Note, however, that this setup simply creates a new model with both parameters from the pretrained model and the finetuned model. In our approach, we dynamically change which components are grafted at each token position (see Figure˜1(b)), allowing us to localize model behavior to both model components and token positions. Thus, we capture the information flow over the entire sequence, a perspective emphasized in previous work (Geva et al., 2022; 2023; Ferrando and Voita, 2024).
Dynamic Weight Grafting
Since our goal is to directly understand the model’s mechanisms, the most direct approach is to just graft in portions of the finetuned model, identifying a subset of the weights and token positions which are sufficient to obtain full finetuned performance. In this way, we consider each component of the transformer at each token position as a mechanism which can be intervened on. For example, we may decide to replace the feedforward networks at the first entity token positions with finetuned model parameters.
@@
More formally, given two models $\bm{\theta}^{A}$ and $\bm{\theta}^{B}$, consider the ordered sequence of their weight matrices $\bm{\theta}^{A}=[\bm{\theta_{1}}^{A}\ldots\bm{\theta_{M}}^{A}]$ and $\bm{\theta}^{B}=[\bm{\theta_{1}}^{B}\ldots\bm{\theta_{M}}^{B}]$. Let $\bm{\gamma}$ be a $1\times M$ mask over these components. Each $\bm{\theta_{c}}^{A}$ corresponds to a specfic model component (e.g. the $W^{Q}$ matrix at the 12th layer, etc.). We define Dynamic Weight Grafting as token-wise, component-wise weight grafting:
$\bm{\tilde{\theta}}_{m}(t)=\begin{cases}\bm{\theta}_{c}^{\text{A}}&\text{if }\gamma_{c}(t)=0\\ \bm{\theta}_{c}^{\text{B}}&\text{if }\gamma_{c}(t)=1\end{cases}$
@@
where $c$ refers to the $c^{\text{th}}$ component of each model and $t$ refes to the token position. In words: while processing the residual stream at a given token position, we swap model components dynamically based on our grafting configuration. This is illustrated in Figure˜1(b). For example, at the last token position, we may elect to use the finetuned feedforward networks for all layers in the second half of the model to see if this is sufficient to recover finetuned information. In our experiments, we consider $\theta_{i}^{\text{A}}$ to be the pretrained model $\theta_{i}^{\text{PRE}}$ and we consider $\theta_{i}^{\text{B}}$ to refer to the finetuned model $\theta_{i}^{\text{SFT}}$.
@@ 3 Experiments & Results
We use four pretrained Transformer-based decoder-only language models in our experiments: Llama3 Grattafiori et al. (2024), Pythia 2.8b Biderman et al. (2023), GPT2-XL Radford et al. (2019), and Gemma 1.1 Gemma Team et al. (2024). Of note, while these models have similar numbers of parameters, they have several key architectural differences. See Table˜2 in Section˜A.3.1 for a comparison of models.
@@
We follow Allen-Zhu and Li (2023) and use templated supervised finetuning data to control the relationship information that models are exposed to during finetuning. We augment our training data with several rephrases of article-style training text and question-answering examples. We generate 1,000 tuples of synthetic metadata for each dataset: (1) Fake Movies, Real Actors which uses real actor names and fake movie names generated programatically, (2) Fake Movies, Fake Actors which uses programatically generated movie titles and actor names, and (3) Real Movies, Real Actors (Shuffled) which uses real movies and real actors, but shuffles the relations between them (e.g. “Keanu Reeves starts in The Departed alongside Meryl Streep”). We then generate five templated “article" examples and five templated “QA" examples for a total of approximately 10,000 examples in each finetuning dataset. See Section˜A.2.2 and Section˜A.2.3 for examples of templates and training examples, Section˜A.2.1 for more detail on dataset creation, and Section˜A.3.3 for training details. In the main body of the paper, we present results for the Fake Movies, Real Actors dataset; results for all datasets are in Appendix˜C. Results are similar across datasets.
@@
Table 1: Relation prompt templates used to test model relation completion capabilities, with examples
@@
Headline {first_actor} {relation} {relation_preposition} a movie {preposition} Brad Pitt starred in a movie with
@@
QA Q: Who {relation} {relation_preposition} a movie {preposition} {first_actor}? A: An actor named Q: Who starred in a movie alongside Brad Pitt? A: An actor named
3.1 Which positions are sufficient for relation completion?